# Communication-Efficient Stochastic Distributed Learning

### v1

@@ I Introduction
In specific terms, the objective in distributed learning is to train a model (e.g., a neural network) with parameters $x\in\mathbb{R}^{n}$ cooperatively across a network of $N$ agents. Each agent has access to a local dataset which defines the local cost as
@@
$\min_{x_{i}\in\mathbb{R}^{n},\ i\in\mathcal{V}}\ \ \frac{1}{N}\sum_{i=1}^{N}f_{i}({x}_{i})\quad\text{s.t.}\ \ x_{1}=x_{2}=\cdots=x_{N}\,,$ (2)
@@
TABLE I: Comparison with the state of the art in stochastic distributed optimization.
@@
K-GT [9] $\tau\div 1$ $2$ $2|\mathcal{N}_{i}|$ $1$ n.c. sub-linear, $\propto\sigma^{2}$
LED [10] $\tau\div 1$ $2$ $|\mathcal{N}_{i}|$ $1$ n.c. s.c. sub-linear, $\propto\sigma^{2}$ linear, $\propto\sigma^{2}$
RandCom [11] $\left\lceil\frac{1}{p}\right\rceil\div 1$ (in mean) $2$ $|\mathcal{N}_{i}|$ $1$ n.c. s.c. sub-linear, $\propto\sigma^{2}$ linear, $\propto\sigma^{2}$
VR-EXTRA/DIGing [12], GT-VR [13] $1\div 1$ $3$ $2|\mathcal{N}_{i}|$ $|\mathcal{B}|$, $m_{i}$ every $\left\lceil\frac{1}{p}\right\rceil$ n.c. s.c. sub-linear, $\to 0$ linear, $\to 0$
GT-SAGA [14, 15] $1\div 1$ $3$ $2|\mathcal{N}_{i}|$ $1$ n.c. s.c. sub-linear, $\to 0$ linear, $\to 0$
GT-SARAH [16] $1\div 1$ $3$ $2|\mathcal{N}_{i}|$ $|\mathcal{B}|$, $m_{i}$ every $\tau$ n.c. sub-linear, $\to 0$
GT-SVRG [14] $1\div 1$ $3$ $2|\mathcal{N}_{i}|$ $1$, $m_{i}$ every $\tau$ s.c. linear, $\to 0$
@@
To effectively tackle this problem, especially when dealing with large datasets that involve sensitive information, distributed methods have become increasingly important. These techniques offer significant robustness advantages over federated learning algorithms [6], as they do not rely on a central coordinator and thus, for example, have not a single point of failure. In particular, both distributed gradient-based algorithms [17, 18, 19, 20], and distributed Alternating Direction Method of Multipliers (ADMM) [21, 22, 23, 24] have proven to be effective strategies for solving such problems.
@@
I-A Comparison with the State of the Art
We compare our proposed algorithms – LT-ADMM and LT-ADMM-VR– with the state of the art. The comparison is holistically summarized in Table I.
Decentralized learning algorithms, as first highlighted in the seminal paper [26] on federated learning, face the fundamental challenge of high communication costs. The authors of [26] address this challenge by designing a communication-efficient algorithms which allows the agents to perform multiple local training steps before each round of communication with the coordinator. However, the accuracy of the algorithm in [26] degrades significantly when the agents have heterogeneous data. Since then, alternative federated learning algorithms, e.g., [27, 28, 29, 30], have been designed to employ local training without compromising accuracy. The interest for communication-efficient algorithms has more recently extended to the distributed set-up, where agents rely on peer-to-peer communications rather than on a coordinator, as in federated learning. Distributed algorithms with local training have been proposed in [31, 9, 10, 11]. In particular, [31, 9, 10] present gradient tracking methods which allow each agent to perform a fixed number of local updates between each communication round. The algorithm in [11], which builds on [28], instead triggers communication rounds according to a given probability distribution, resulting in a time-varying number of local training steps. Another related algorithm is that of [32], which allows for both multiple consensus and gradient steps in each iteration. However, this algorithm requires a monotonically increasing number of communication rounds in order to guarantee exact convergence. A stochastic version of [32] was then studied in [33]. The algorithm has inexact gradient evaluations, but only allows for multiple consensus steps.
@@ I-A Comparison with the State of the Art
We conclude by providing a summary of the features of the algorithms compared so far in Table I. First, we classify them based on whether they use or not variance reduction and local training. For the latter, we report the ratio of gradient steps to communication rounds that characterizes each algorithm, with a ratio of $1\div 1$ signifying that no local training is used. Notice that LT-ADMM-VR is the only algorithm to use both variance reduction and local training, while the other only use one technique. We then compare the number of variables stored by the agents when they deploy each algorithm (disregarding temporary variables). We notice that the variable storage of LT-ADMM and LT-ADMM-VR, differently from the alternatives, scales with the size of an agent’s neighborhood; this is due to the use of distributed ADMM as the foundation of our proposed algorithms [21]. We see that [10, 11], LT-ADMM, and LT-ADMM-VR require one communication per neighbor, while the other methods require two communications per neighbor. We also compare the algorithms by the computational complexity of the gradient estimators they employ, namely, the number of component gradient evaluations needed per local training iteration. The algorithms of [9, 10, 11, 14] use a single data point to estimate the gradient, while [12, 16], LT-ADMM, LT-ADMM-VR can apply mini-batch estimators that use a subset $\mathcal{B}$ of the local data points. The use of mini-batches yields more precise gradient estimates and increased flexibility. However, we remark that the gradient estimators used in [12, 16, 14], LT-ADMM, LT-ADMM-VR require a registry of component gradient evaluations, which needs to be refreshed entirely at fixed intervals. This coincides with the evaluation of a full gradient, and thus requires $m_{i}$ component gradient evaluations. Finally, we compare the algorithms’ convergence. We notice that all algorithms, except for [14], provide (sub-linear) convergence guarantees for convex and nonconvex problems. Additionally, some works show linear convergence for strongly convex problems. We further distinguish between algorithms which achieve exact convergence due to the use of variance reduction, or inexact convergence with an error proportional to the stochastic gradient variance ($\propto\sigma^{2}$).
@@
The outline of the paper is as follows. Section II formulates the problem at hand, and presents the proposed algorithms design. Section III analyzes their convergence, and discusses the results. Section IV reports and discusses numerical results comparing the proposed algorithms with the state of the art. Section V presents some concluding remarks.
@@
[…]
II Problem Formulation and Algorithm Design
@@
II-A Problem Formulation
We target the solution of (2) over a (undirected) graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{1,...,N\}$ is the set of $N$ agents, and $\mathcal{E}\subset\mathcal{V}\times\mathcal{V}$ is the set of edges $(i,j)$, $i,j\in\mathcal{V}$. In particular, we assume that the local costs $f_{i}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ are in the empirical risk minimization form (1). We make the following assumptions for (2), which commonly underpin the convergence analysis of distributed algorithms.
Assumption 1.
@@ II-A Problem Formulation
When, in the following, we specialize our results to convex scenarios, we resort to the further assumption below.
Assumption 3.
@@ II-B Algorithm Design
$z_{ij,k+1}$ $=\frac{1}{2}\left(z_{ij,k}-z_{ji,k}-2\rho x_{j,k+1}\right)\,,$ (3b)
@@
$\begin{split}&\phi_{i,k}^{0}=x_{i,k}\,,\\ &\phi_{i,k}^{t+1}=\phi_{i}^{t}-\gamma\Bigg{(}g_{i}(\phi_{i,k}^{t})+\rho\left|\mathcal{N}_{i}\right|\phi_{i,k}^{t}-\sum_{j\in\mathcal{N}_{i}}z_{ij,k}\Bigg{)},\\ &\qquad t=0,\ldots,\tau-1\,,\\ &x_{i,k+1}=\phi_{i,k}^{\tau}\,,\end{split}$ (4)
where $g_{i}(\phi_{i,k}^{\ell})$ is an estimate of the gradient $\nabla f_{i}$. The resulting algorithm is a distributed gradient method, with the difference that each communication round (3b) is preceded by $\tau>1$ local gradient evaluations. This is an application of the local training paradigm [10]. We remark that the convergence of the proposed algorithm rests on the initialization $\phi_{i,k}^{0}=x_{i,k}$, which enacts a feedback loop on the local training. In general, without this initialization, exact convergence cannot be achieved [22].
The local training (4) requires a local gradient evaluation or at least its estimate. In the following, we introduce two different estimator options. Notice that the gradient of the penalty term, $\rho\left|\mathcal{N}_{i}\right|\phi_{i,k}^{t}-\sum_{j\in\mathcal{N}_{i}}z_{ij,k}$, is exactly known and does not need an estimator. The most straightforward idea is to simply employ a local gradient $g_{i}(\phi)=\nabla f_{i}(\phi)$. However, in learning applications, the agents may store large datasets ($m_{i}\gg 1$). Therefore, computing $\nabla f_{i}(\phi)$ becomes computationally expensive. To remedy this, the agents can instead use stochastic gradients, choosing
@@
where $\mathcal{B}_{i}$ are randomly drawn indices from $\{1,\ldots,m_{i}\}$, with $|\mathcal{B}_{i}|<m_{i}$. While reducing the computational complexity of the local training iterations, the use of stochastic gradients results in inexact convergence. The idea therefore is to employ a gradient estimator based on a variance reduction scheme. In particular, we use the scheme proposed in [25], which demonstrates faster convergence in practice and is characterized by the following procedure. Each agent maintains a table of component gradients $\{\nabla f_{i,h}(r_{i,h,k}^{t})\},h=1,\ldots,m_{i}$, where $r_{i,h,k}^{t}$ is the most recent iterate at which the component gradient was evaluated. This table is reset at the beginning of every new local training (that is, for any $k\in\mathbb{N}$ when $t=0$). Using the table, the agents then estimate their local gradients as
@@
The resulting algorithms are reported in Algorithm 1.
@@
1:For each node $i$, initialize $x_{i,0}=z_{ij,0}$, $j\in\mathcal{N}_{i}$. Set the penalty $\rho$, the number of local training steps $\tau$, and the local step-size $\gamma$.
2:for $k=0,1,\ldots$ every agent $i$ do
@@
III Convergence Analysis and Discussion
@@
where $F(x)=\frac{1}{N}\sum_{i=1}^{N}f_{i}(x)$ and $\bar{x}_{k}=\frac{1}{N}\sum_{i=1}^{N}x_{i}$. If the agents’ converge to a stationary point of (2), then $\mathcal{D}_{k}\to 0$.
III-A Convergence with SGD
We start by characterizing the convergence of Algorithm 1 when the agents use SGD during local training. To this end, we make the following standard assumption on the variance of the gradient estimators.
Assumption 4.
@@ III-A Convergence with SGD
We are now ready to state our convergence results. All the proofs are deferred to the Appendix.
Theorem 1 (Nonconvex case).
Let Assumptions 1, 2, and 4 hold. If the local step-size $\gamma$ is chosen sufficiently small (see (10) for the precise bound), then the output of LT-ADMM satisfies:
$\frac{1}{K}\sum_{k=0}^{K-1}\mathcal{D}_{k}\leq\mathcal{O}\left(\frac{F(\bar{x}_{0})-F(x^{*})}{K\gamma\tau}\right)+\mathcal{O}\left(\gamma\tau\sigma^{2}\right)+\mathcal{O}\left(\frac{\|\widehat{\mathbf{d}}_{0}\|^{2}}{\rho KN}\right),$ (8)
where $x^{*}$ is a stationary point of (2), and $\|\widehat{\mathbf{d}}_{0}\|$ is related to the initial conditions (see (26)).
@@
Corollary 1 (Convex case).
@@
Remark 1 (Exact convergence).
@@
III-B Convergence with Variance Reduction
The results of the previous sections show that only inexact convergence can be achieved when employing SGD. The following results show how Algorithm 1 achieves exact convergence when using variance reduction.
Theorem 2 (Nonconvex case).
Let Assumptions 1, 2, and 4 hold. If the local step-size $\gamma$ is chosen sufficiently small (see (11) for the precise bound), then the output of LT-ADMM converges to a stationary point $x^{*}$ of (2), and in particular it holds:
$\frac{1}{K}\sum_{k=0}^{K-1}\mathcal{D}_{k}\leq\mathcal{O}\left(\frac{F\left(\bar{x}_{0}\right)-F(x^{*})}{K\gamma\tau}\right)+\mathcal{O}\left(\frac{\|\widehat{\mathbf{d}}_{0}\|^{2}}{\rho^{2}KN}\right)$ (9)
Corollary 2 (Convex case).

### v2

@@ 1 Introduction
In specific terms, the objective in distributed learning is to train a model (e.g., a neural network) with parameters $x\in\mathbb{R}^{n}$ cooperatively across a network of $N$ agents. Each agent $i$ has access to a local dataset which defines the local cost as
@@
$\min_{\begin{subarray}{c}x_{i}\in\mathbb{R}^{n}\\ i={1,...N}\end{subarray}}\ \ \frac{1}{N}\sum_{i=1}^{N}f_{i}({x}_{i})\quad\text{s.t.}\ \ x_{1}=x_{2}=\cdots=x_{N}\,$,} (2)
@@
Table 1: Comparison with the state of the art in stochastic distributed optimization.
@@
K-GT [9] $\tau\div 1$ 22 $2|\mathcal{N}_{i}|$ 11 n.c. sub-linear, $\propto\sigma^{2}$
LED [10] $\tau\div 1$ 22 $|\mathcal{N}_{i}|$ 11 n.c. s.c. sub-linear, $\propto\sigma^{2}$ linear, $\propto\sigma^{2}$
RandCom [11] $\left\lceil\frac{1}{p}\right\rceil\div 1$ (in mean) 22 $|\mathcal{N}_{i}|$ 11 n.c. s.c. sub-linear, $\propto\sigma^{2}$ linear, $\propto\sigma^{2}$
VR-EXTRA/DIGing [12], GT-VR [13] $1\div 1$ 33 $2|\mathcal{N}_{i}|$ $|\mathcal{B}|$, $m_{i}$ every $\left\lceil\frac{1}{p}\right\rceil$ n.c. s.c. sub-linear, $\to 0$ linear, $\to 0$
GT-SAGA [14, 15] $1\div 1$ 33 $2|\mathcal{N}_{i}|$ 11 s.c. n.c. sub-linear, $\to 0$ linear, $\to 0$
GT-SARAH [16] $1\div 1$ 33 $2|\mathcal{N}_{i}|$ $|\mathcal{B}|$, $m_{i}$ every $\tau$ n.c. sub-linear, $\to 0$
GT-SVRG [14] $1\div 1$ 33 $2|\mathcal{N}_{i}|$ 11, $m_{i}$ every $\tau$ s.c. linear, $\to 0$
@@
To effectively tackle this problem, especially when dealing with large datasets that involve sensitive information, distributed methods have become increasingly important. These techniques offer significant robustness advantages over federated learning algorithms [6], as they do not rely on a central coordinator and thus, for example, have not a single point of failure. In particular, both distributed gradient-based algorithms [17, 18, 19, 20], and distributed Alternating Direction Method of Multipliers (ADMM) [21, 22, 23, 24] have proven to be effective strategies for solving such problems. ADMM-based algorithms have demonstrated strong robustness to practical constraints (see e.g. [22] and references therein), although this often comes at the cost of higher computational complexity compared to gradient-based methods. In this work, we propose novel ADMM-based algorithms that retain the computational efficiency characteristic of gradient methods.
@@
1.1 Comparison with the state of the art
We compare our proposed algorithms – LT-ADMM and LT-ADMM-VR– with the state of the art. The comparison is holistically summarized in Table 1.
Decentralized learning algorithms, as first highlighted in the seminal paper [26] on federated learning, face the fundamental challenge of high communication costs. The authors of [26] address this challenge by designing a communication-efficient algorithms which allows the agents to perform multiple local training steps before each round of communication with the coordinator. However, the accuracy of the algorithm in [26] degrades significantly when the agents have heterogeneous data. Since then, alternative federated learning algorithms, e.g., [27, 28, 29, 30], have been designed to employ local training without compromising accuracy. The interest for communication-efficient algorithms has more recently extended to the distributed set-up, where agents rely on peer-to-peer communications rather than on a coordinator as in federated learning. Distributed algorithms with local training have been proposed in [31, 9, 10, 11]. In particular, [31, 9, 10] present gradient tracking methods which allow each agent to perform a fixed number of local updates between each communication round. The algorithm in [11], which builds on [28], instead triggers communication rounds according to a given probability distribution, resulting in a time-varying number of local training steps. Another related algorithm is that of [32], which allows for both multiple consensus and gradient steps in each iteration. However, this algorithm requires a monotonically increasing number of communication rounds in order to guarantee exact convergence. A stochastic version of [32] was then studied in [33]. The algorithm has inexact gradient evaluations, but only allows for multiple consensus steps. An alternative approach to reducing the frequency of communications is to employ event-triggering, see e.g. [34], where messages are exchanged only when a certain condition is met.
@@ 1.1 Comparison with the state of the art
We conclude by providing in Table 1 a summary of the key features of the algorithms discussed above. This table focuses on methods that employ the mechanisms of primary interest in this work – local training and variance reduction. First, we classify them based on whether they use or not variance reduction and local training. For the latter, we report the ratio of gradient steps to communication rounds that characterizes each algorithm, with a ratio of $1\div 1$ signifying that no local training is used. Notice that LT-ADMM-VR is the only algorithm to use both variance reduction and local training, while the other only use one technique. We then compare the number of variables stored by the agents when they deploy each algorithm (disregarding temporary variables). We notice that the variable storage of LT-ADMM and LT-ADMM-VR, differently from the alternatives, scales with the size of an agent’s neighborhood; this is due to the use of distributed ADMM as the foundation of our proposed algorithms [21]. We see that [10, 11], LT-ADMM, and LT-ADMM-VR require one communication per neighbor, while the other methods require two communications per neighbor. We also compare the algorithms by the computational complexity of the gradient estimators they employ, namely, the number of component gradient evaluations needed per local training iteration. The algorithms of [9, 10, 11, 14] use a single data point to estimate the gradient, while [12, 16], LT-ADMM, LT-ADMM-VR can apply mini-batch estimators that use a subset $\mathcal{B}$ of the local data points. The use of mini-batches yields more precise gradient estimates and increased flexibility. However, we remark that the gradient estimators used in [12, 16, 14], LT-ADMM, LT-ADMM-VR require a registry of component gradient evaluations, which needs to be refreshed entirely at fixed intervals. This coincides with the evaluation of a full gradient, and thus requires $m_{i}$ component gradient evaluations. Finally, we compare the algorithms’ convergence. We notice that all algorithms, except for [14], provide (sub-linear) convergence guarantees for convex and nonconvex problems. Additionally, some works show linear convergence for strongly convex problems. We further distinguish between algorithms which achieve exact convergence due to the use of variance reduction, or inexact convergence with an error proportional to the stochastic gradient variance ($\propto\sigma^{2}$).
@@
The outline of the paper is as follows. Section 2 formulates the problem at hand, and presents the proposed algorithms design. Section 3 analyzes their convergence, and discusses the results. Section 4 reports and discusses numerical results comparing the proposed algorithms with the state of the art. Section 5 presents some concluding remarks.
@@
$\nabla f$ denotes the gradient of a differentiable function $f$. Given a matrix $A\in\mathbb{R}^{n\times n}$, $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ denotes the smallest and largest eigenvalue of $A$, respectively. $A>0$ represents that matrix $A$ is positive definite. With $n\in\mathbb{N}$, we let $\mathbf{1}_{n}\in\mathbb{R}^{n}$ be the vector with all elements equal to $1$, $\mathbf{I}\in\mathbb{R}^{n\times n}$ the identity matrix and $\mathbf{0}\in\mathbb{R}^{n\times n}$ the zero matrix. $\langle x,y\rangle=\sum_{h=1}^{n}x_{h}y_{h}$ represents the standard inner product of two vectors $x,y\in\mathbb{R}^{n}$. $\|\cdot\|$ denotes the Euclidean norm of a vector and the matrix-induced 2-norm of a matrix. The proximal of a cost $f$, with penalty $\rho>0$, is defined as $\operatorname{prox}_{f}^{\rho}(y)=\operatorname*{arg\,min}_{y\in\mathbb{R}^{n}}\left\{f(y)+\frac{1}{2\rho}\left\lVert y-x\right\rVert^{2}\right\}$.
2 Problem Formulation and Algorithm Design
@@
2.1 Problem formulation
We target the solution of (2) over a (undirected) graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{1,...,N\}$ is the set of $N$ agents, and $\mathcal{E}\subset\mathcal{V}\times\mathcal{V}$ is the set of edges $(i,j)$, $i,j\in\mathcal{V}$. In particular, we assume that the local costs $f_{i}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ are in the empirical risk minimization form (1). We make the following assumptions for (2), which are commonly used to support the convergence analysis of distributed learning algorithms (see e.g. [9, 10, 11, 12, 15, 16]).
Assumption 1
@@ 2.1 Problem formulation
When, in the following, we specialize our results to convex scenarios, we resort to the additional assumption below.
Assumption 3
@@ 2.2 Algorithm design
$z_{ij,k+1}$ $=\frac{1}{2}\left(z_{ij,k}-z_{ji,k}+2\rho x_{j,k+1}\right)\,$, (3b)
@@
$\phi_{i,k}^{0}=x_{i,k}\,$,
$\phi_{i,k}^{t+1}=\phi_{i}^{t}-\Bigg(\gamma g_{i}(\phi_{i,k}^{t})+\beta\Big(\rho\left|\mathcal{N}_{i}\right|x_{i,k}-\sum_{j\in\mathcal{N}_{i}}z_{ij,k}\Big)\Bigg),$
$\qquad t=0,\ldots,\tau-1\,$,
$x_{i,k+1}=\phi_{i,k}^{\tau}\,$, (4)
where $\gamma$, $\beta$ are the positive step-sizes, and $g_{i}(\phi_{i,k}^{\ell})$ is an estimate of the gradient $\nabla f_{i}$. Notice that for efficiency’s sake we “freeze” the penalty term $\rho\left|\mathcal{N}_{i}\right|x_{i,k}-\sum_{j\in\mathcal{N}_{i}}z_{ij,k}$, and for flexibility we multiply gradient estimate and penalty term by two different step-sizes. The resulting algorithm is a distributed gradient method, with the difference that each communication round (3b) is preceded by $\tau>1$ local gradient evaluations. This is an application of the local training paradigm [10]. We remark that the convergence of the proposed algorithm rests on the initialization $\phi_{i,k}^{0}=x_{i,k}$, which enacts a feedback loop on the local training. In general, without this initialization, exact convergence cannot be achieved [22].
The local training (4) requires a local gradient evaluation or at least its estimate. In the following, we introduce two different estimator options. Notice that the gradient of the penalty term, $\rho\left|\mathcal{N}_{i}\right|x_{i,k}-\sum_{j\in\mathcal{N}_{i}}z_{ij,k}$, is exactly known (and frozen) and does not need an estimator. The most straightforward idea is to simply employ a local gradient $g_{i}(\phi)=\nabla f_{i}(\phi)$. However, in learning applications, the agents may store large datasets ($m_{i}\gg 1$). Therefore, computing $\nabla f_{i}(\phi)$ becomes computationally expensive. To remedy this, the agents can instead use stochastic gradients, choosing
@@
where $\mathcal{B}_{i}$ are randomly drawn indices from $\{1,\ldots,m_{i}\}$, with $|\mathcal{B}_{i}|<m_{i}$. While reducing the computational complexity of the local training iterations, the use of stochastic gradients results in inexact convergence. The idea, therefore, is to employ a gradient estimator based on a variance reduction scheme. In particular, we adopt the scheme proposed in [25], characterized by the following procedure. Each agent maintains a table of component gradients $\{\nabla f_{i,h}(r_{i,h,k}^{t})\},h=1,\ldots,m_{i}$, where $r_{i,h,k}^{t}$ is the most recent iterate at which the component gradient was evaluated. This table is reset at the beginning of every new local training (that is, for any $k\in\mathbb{N}$ when $t=0$). Using the table, the agents then estimate their local gradients as
@@
Selecting the stochastic gradient estimator (5) yields the proposed algorithm LT-ADMM, while selecting the variance reduction scheme (6) yields the proposed algorithm LT-ADMM-VR. The two methods are reported in Algorithm 1.
@@
1:For each node $i$, initialize $x_{i,0}=z_{ij,0}$, $j\in\mathcal{N}_{i}$. Set the penalty $\rho$, the number of local training steps $\tau$, the number of iterations $K$, and the local step-size $\gamma$, $\beta$.
2:for $k=0,1,\ldots,K-1$ every agent $i$ do
@@
3 Convergence Analysis and Discussion
@@
where $F(x)=\frac{1}{N}\sum_{i=1}^{N}f_{i}(x)$ and $\bar{x}_{k}=\frac{1}{N}\sum_{i=1}^{N}x_{i,k}$. If the agents converge to a stationary point of (2), then $\mathcal{D}_{k}\to 0$. We note that this performance measure is standard in the literature on stochastic gradient methods and distributed optimization [10, 11]. Although it does not, in general, imply almost sure convergence of the sequence $\{\mathcal{D}_{k}\}_{k=1}^{K}$, it provides meaningful performance guarantees in expectation. Specifically, if the index $K^{\prime}$ is selected uniformly at random from $1,\ldots,K$}, then $\mathbb{E}[\mathcal{D}_{K^{\prime}}]=\frac{1}{K}\sum_{k=0}^{K-1}\mathcal{D}_{k}$.
3.1 Convergence with SGD
We start by characterizing the convergence of Algorithm 1 when the agents use SGD during local training (LT-ADMM). To this end, we make the following standard assumption on the variance of the gradient estimators, see e.g. [9, 10].
Assumption 4
@@ 3.1 Convergence with SGD
We are now ready to state our convergence results. All the proofs are deferred to the Appendix, where Appendix 6 provides a sketch of the proofs, followed by the full proofs. We remark that prior analyses of distributed ADMM based on operator-theoretic approaches [21, 22] are not directly applicable to LT-ADMM, and the convergence proofs must therefore be specifically tailored to this algorithm.
Theorem 1 (Nonconvex case)
Let Assumptions 1, 2, and 4 hold. If the local step-sizes satisfy $\gamma\leq\mathcal{O}(\frac{\lambda_{l}}{L\tau^{2}})<\gamma_{\text{sgd}}\coloneqq\min_{i=1,2,\ldots,6}\bar{\gamma}_{i}$ (see (10) in Appendix 7.1 for the precise bound), and $1/(\tau\lambda_{u}\rho)\leq\beta<2/(\tau\lambda_{u}\rho)$, then the output of LT-ADMM satisfies:
$\frac{1}{K}\sum_{k=0}^{K-1}\mathcal{D}_{k}\leq\mathcal{O}\left(\frac{F(\bar{x}_{0})-F(x^{*})}{K\gamma\tau}\right)+\mathcal{O}\left(\gamma\tau\sigma^{2}\right)+{\mathcal{O}\left(\frac{\|\widehat{\mathbf{d}}_{0}\|^{2}}{\rho^{2}KN}\right)},$ (8)
where $x^{*}$ is a stationary point of (2), $\lambda_{u}$ is the largest eigenvalue of the graph $\mathcal{G}$’s Laplacian matrix, $\lambda_{l}$ is the smallest nonzero eigenvalue of graph $\mathcal{G}$’ Laplacian matrix, and $\|\widehat{\mathbf{d}}_{0}\|$ is related to the initial conditions (see (26)).
@@
Corollary 1 (Convex case)
@@
Remark 1 (Exact convergence)
@@
3.2 Convergence with variance reduction
The results of the previous section shows that only inexact convergence can be achieved when employing SGD. The following results show how Algorithm 1 achieves exact convergence when using variance reduction (LT-ADMM-VR).
Theorem 2 (Nonconvex case)
Let Assumptions 1, 2 hold. If the local step-sizes satisfy $\gamma\leq\mathcal{O}(\frac{\lambda_{l}}{L\tau^{3}})<\gamma_{\text{vr}}\coloneqq\min_{i=1,7,8,\ldots,15}\bar{\gamma}_{i}$ (see (11) in Appendix 7.1 for the precise bound), and $1/(\tau\lambda_{u}\rho)\leq\beta<2/(\tau\lambda_{u}\rho)$, then the output of LT-ADMM-VR converges to a stationary point $x^{*}$ of (2), and in particular it holds:
$\frac{1}{K}\sum_{k=0}^{K-1}\mathcal{D}_{k}\leq\mathcal{O}\left(\frac{F\left(\bar{x}_{0}\right)-F(x^{*})}{K\gamma\tau}\right)+{\mathcal{O}\left(\frac{\|\widehat{\mathbf{d}}_{0}\|^{2}}{\rho^{2}K}\right),}$ (9)
where $x^{*}$ is a stationary point of (2), $\lambda_{u}$ is the largest eigenvalue of the graph $\mathcal{G}$’s Laplacian matrix, $\lambda_{l}$ is the smallest nonzero eigenvalue of graph $\mathcal{G}$’ Laplacian matrix, and $\|\widehat{\mathbf{d}}_{0}\|$ is related to the initial conditions (see (26)).
Corollary 2 (Convex case)